# SynthBA: Reliable Brain Age Estimation Across Multiple MRI Sequences and Resolutions

### v1

@@ III-D Clinical evaluation
In this experiment, we conduct a concise clinical evaluation of SynthBA, examining the relationship between the ADAS-Cog13 score and brain PAD determined by the SynthBA model. The ADAS-Cog13 [26] score serves as a cognitive evaluation tool specifically designed for assessing cognitive function in AD patients. Using the SynthBA-g model, we extract predicted brain age data from subjects within the AD-EVAL dataset. Brain PAD is computed as the difference between the predicted brain age and the chronological age. Our analysis reveals a significant Pearson correlation of 0.306 ($p<0.001$) between brain PAD and the ADAS-Cog-13 score among AD subjects, as depicted in Figure 3. Notably, Brain PADs computed via SynthBA effectively replicate key associations linking brain PAD with cognitive decline [27].

### v2

@@ III-D Clinical evaluation
In this experiment, we conduct a clinical evaluation of SynthBA, examining the relationship between the ADAS-Cog13 score and brain PAD determined by the SynthBA model. The ADAS-Cog13 [27] score serves as a cognitive evaluation tool specifically designed for assessing cognitive function in AD patients. Using the best SynthBA configuration (SynthBA-g), we extract predicted brain age data from subjects within the AD-EVAL dataset and compute the brain PAD. Our analysis reveals a significant Pearson correlation of 0.275 ($p<0.001$) between brain PAD and the ADAS-Cog-13 score among AD subjects, as depicted in Figure 3.